Evaluate the infinite geometric series. Round to the nearest hundredth if necessary.
50 + 10 + 2 + ...

Answer by jim_thompson5910(35256)

50 + 10 + 2 + ... = S

S = a/(1-r)

S = 50/(1-1/5)

S = 50/(4/5)

S = 50*(5/4)

S = 250/4

S = 125/2

S = 62.5

50 + 10 + 2 + ... = S

50 + 10 + 2 + ... = 62.5